# Pay Less Attention to Function Words for Free Robustness of Vision-Language Models

### v2

@@ 4.1 Main Results
T2IR/I2TR. Results of T2IR/I2TR on ALBEF, TCL, and BLIP for all methods are given in Table.2. Overall on Flickr, our FDA consistently exhibits the best robustness with the best clean performance on all models over all other baselines, yielding a 22.26/14.69%, 14.29/13.55%, and 51.60/56.36% average ASR drop over all 3 attacks on ALBEF/TCL/BLIP for $\epsilon=2/4$, with a negligible 0.30/0.10%, 0.50/0.22%, and 0.70/46% performance drops in R@1 for T2IR/I2TR on each model, respectively. On MSCOCO, similar patterns exist as our FDA boosts the ASR drop by 9/14% for $\epsilon=2/4$ with a 0.1% clean performance boost.
i. Attack-wise, on Flickr, our FDA exhibits the best defense against PGD and MAPGD in 22 out of 24 results, leading TeCoA/FARE by 60/65% on the BLIP model, demonstrating the effectiveness of FDA in enhancing robustness against various attacks. For the strongest adaptive attack, MAPGD, our FDA maintains its lead over TeCoA and FARE on ABLEF and BLIP, with an average lead by over 10%. Although our FDA shows more vulnerability against APGD on the TCL model, it retains the best comprehensive robustness of the other two baselines, yielding a 10-20% lead for $\epsilon=2/4$. It is noticeable that TeCoA/FARE becomes ineffective for all attacks with $\epsilon=4$, while our FDA retains its effectiveness facing stronger attacks. Similar trends also exist on MSCOCO.
ii. Performance-wise, all baseline methods suffered from a performance drop by an average4/3/9% on ALBEF/TCL/BLIP. Nevertheless, our FDA only causes minor or little drops of less than 1% for all models, yielding a lead of TeCoA and FARE by approximately 4%, 3%, and %/7 on average, demonstrating the feasibility of paying less attention to function words for free robustness.
@@
Table 4: Robustness evaluations on ALBEF using FDA as a plug-and-play tool with TeCoA and FARE against targeted and untargeted attacks for Text-to-Image/Image-to-Text Retrieval. Results are averaged over T2IR and I2TR. Full results are provided in Sec.C of the Appendix. FDA consistently boosts clean performance and/or robustness against all attacks.
@@
T2IR I2TR PGD APGD MAPGD PGD APGD MAPGD Robustness ($\uparrow$)
@@
Table 5: Robustness evaluations of FDA as a plug-and-play tool with TeCoA and FARE against targeted and untargeted attacks for Visual Grounding. Results are averaged over Test-A and Test-B. Full results are provided in Sec.C of the Appendix. FDA consistently boosts clean performance and robustness against all attacks.
@@
4.2 Untargeted Attacks
Apart from targeted attacks, we further evaluate the robustness against untargeted attacks. Thus, we retrained all models using TeCoA/FARE and their combination with our FDA to validate the effectiveness of FDA in defending against untargeted attacks.
For T2IR/I2TR, as presented in Table.4. Overall, we find FDA consistently boosts the robustness of TeCoA and FARE for all untargeted attacks on all models. Specifically, the scalability of FDA also applies after combining with TeCoA/FARE: both methods benefit more from FDA on the larger backbone of BLIP, yielding a 4/3% robustness boost. Furthermore, we notice that FDA also boosts the clean performance of both methods on ALBEF considerably, besides the improvement in robustness. For VG, we observe identical patterns: implementing FDA yields a solid robustness gain. For example, FARE experiences a significant robustness enhancement regarding untargeted attacks by 5%. In sum, our FDA compatibility works with both TeCoA and FARE to further boost their robustness against untargeted attacks.
@@ 4.3 Ablation Study
We now provide comprehensive studies on untargeted attacks, hyperparameters of FDA: encoder, dictionary, layer/head, and zero-shot performance. (See full results in Sec.D of the Appendix.)
4.3.1 FDA v.s. Masking & Other DA
We start by providing comparisons of our FDA and fine-tuning models by directly masking function words. We also include content words and nouns for thorough evaluation. As for variant of FDA, we further compare our FDA with Adjective DA (ADA) and Determiner DA (DDA). Specifically, we choose determiners (DET) and adjectives because DET indicates using a small subset (i.e., a/an/the) of function words, while ADJ adopts a completely different set of words. Results are presented in Table.6. Note: We only test on PGD and APGD since MAPGD is not applicable for nouns and content words.
Table 6: Comparison bewteen fine-tuning ALBEF by directly removing content words (CONT), nouns (NOUN), function words (FUNC), Determiner DA (DDA), and Adjective DA (ADA). $\Delta_{ASR}$ is presented using the average results for PGD and APGD. De-Attention shows significan advantages over directly masking and de-attentioning other words.
Maksd Clean (R@1) ($\uparrow$) Avg ASR Drop
Words T2IR I2TR $\Delta_{ASR}$ ($\uparrow$)
N/A 95.90 85.60 -
CONT 21.50 11.10 -
NOUN 68.60 44.62 -
FUNC 94.00 80.86 $\uparrow$  1.56
DDA 95.60 85.42 $\uparrow$  9.28
ADA 95.50 85.38 $\uparrow$ 15.10
FDA 95.60 85.50 $\uparrow$ 23.07
4.3.2 FDA v.s. Adaptive Wwords selection
[…]
in Dictionary T2IR I2TR $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ Avg
SIM-N 25.95 95.90 85.50 2.44 12.81 $\uparrow$  7.62
SIM-$2\delta$ 74.53 95.60 85.32 9.38 13.86 $\uparrow$  8.39
SIM-$\delta$ 79.49 95.30 85.38 12.85 11.54 $\uparrow$ 12.41
FDA 100.00 95.90 85.50 27.61 18.53 $\uparrow$ 23.07
@@
Table 9: Ablation studies on the layer/head index $L/H$ of FDA on Text-to-Image/Image-to-Text Retreival on ALBEF, TCL and BLIP. Results are averaged over T2IR/I2TR. Shallower layers/heads (smaller $L/H$) consistently outperform over others on retrieval tasks.
@@
Table 10: Ablation studies on the layer/head index $L/H$ of FDA on Visual Grounding Retrieval on ALBEF, TCL, and BLIP. Results are averaged over Test A/B splits.
@@
First of all, masking content words and nouns yields the largest performance drop, making it unviable for robustness evaluation. This aligns with the intuition that these words carry extensive semantic information crucial for VLM tasks. Furthermore, masking function words leads to evident performance drop ($\sim$3%) and brings negligible robustness ($\sim$1%).As for DDA, as a subset of FDA, it shows almost identical clean performance, with a significant drop in robustness, indicating insufficient de-attentioning. ADA also shows subpar performance compared with FDA. Overall, we find that FDA leads the clean and adversarial performance among other variants.
4.3.3 Hyperparameters
The implementation of FDA, especially the macro-hypaerparameters influencing where to implement, would largely impact the subsequent performance of models. We first provide relative ablation studies to help understand the mechanics and design of our FDA.
Encoders&Dictionary. We start by comparing three implementations: FDA on text encoders, fusion encoders, and both, denoted as $\mathcal{T}$, $\mathcal{H}$, and $\mathcal{T}\&\mathcal{H}$. As presented in the top rows of Table.8, we find that $\mathcal{T}$ performs the worst among all, indicating that an early subtraction is insufficient for removing such subtraction. Although $\mathcal{T}\&\mathcal{H}$ provides a significant robustness boost, it costs an evident 2% performance drop on performance, implying that subtraction on both encoders is too strong and potentially causes contextual distortion. $\mathcal{H}$ performs the best as it helps models concentrate while preserving the contextual meaning.
For the dictionary, we use the off-the-shelf stopwords dictionary in [9], containing 208 words/symbols, denoted as Full Dict. Furthermore, we use a shortlisted dictionary, by only using the most commonly used function words, containing 93 crucial function words, denoted as Shortlisted Dict. Both dictionary settings are trained with FDA $L^{a}ll$ to maximize their impacts on training. As presented in the lower row of Table.8, there are no significant performance gaps between the two settings, with Full Dict performing slightly worse regarding both clean and adversarial examples. We attribute the minor degradation to the length of the stopwords dictionary, which could unnecessarily skim words and distort the context. We provide the shortlisted dictionary in Sec.E of the Appendix.
Attention Head &Layer. We then investigate the index of the layers $L$ and attention heads $H$ for retrieval and grounding, as presented in Table.9 and Table.10. Specifically, we train a series models using FDA but using different $L$ and $H$: for layers, we use all, 0-1, and 0 layers, denoted as $L^{all},L^{0-1},L^{0}$; for attention heads, we use all heads, 1st half (0-5) and the second half (6-11), denoted as $H^{all},H^{0-5},H^{6-11}$. For T2IR/I2TR, as shown in Table.9, we find that the shallow implementations of FDA, i.e., $L^{0}/L^{0-1},H^{0-5}$ consistently yield the best performance on robustness on all models. Specifically, $L^{0},H^{0-5}$ constantly achieves the best clean performance, leading other counterparts by 0.1-0.2%.
We further test the leading 3 settings on retrieval tasks, i.e., $L^{all}/L^{0}/L^{0-1},H^{0-5}$ on VG. As shown in 10, we find the shallow $L^{0-1},H^{0-5}$ settings still top w.r.t. both adversarial and clean examples, leading other settings by 10-20%/0.3-0.4%, respectively.
Overall, while FDA behaves slightly differently across settings/tasks, its effectiveness remains solid and insensitive to the head/layer parameters.
4.3.4 Zero-shot Performance
Finally, we adopt the three settings of FDA without finetuning to evaluate the zero-shot performance on different tasks (T2IR/I2TR/VG) on ALBEF and BLIP ($H^{0-5}$ is omitted and unchanged for all FDA). Results are presented in Table.11. We find that $L^{all}$ performs the best for all tasks and all models. This not only suggests that $L^{all}$ serves as the most generalizable setting for multiple VL tasks and models, but also implies the feasibility of FDA for performance boost on zero-shot tasks.
@@ 4.4 Analysis and Visualization
We notice that the results of APGD and MAPGD somewhat worsen after adversarial finetuning, e.g., TeCoA and FARE on ALBEF, FARE on BLIP in Table.2, etc. As previously illustrated in Fig.2, defending against targeted attacks requires a more aligned vision-language embedding. Consequently, we hypothesize that such abnormality potentially originates from the disruption in vision-language alignment brought by adversarial noise for enhanced robustness.
Figure 3: Left: T-SNE of the vision-language embedding of vanilla VLM, FDA, FARE, and TeCoA. Our FDA is the most aligned model. Right: Comparison of text-image similarity for vanilla VLM versus VLM + FDA. Our FDA yields better alignment with larger similarities and smaller variances.
@@
4.5 Limitaiton
Besides subtraction, FDA could be potentially improved through a modular or algorithmic approach for more refined removal. Furthermore, we did not implement FDA to fine-tune a larger VLM or verify the effectiveness of FDA using LoRa due to the hardware limitation. Finally, our FDA is designed for backbones with a fusion encoder and thus not directly implementable for CLIP and other similar backbones. However, we believe implementation on CLIP-like models would be a valuable exploration for future work.
@@ 5 Conclusion
We propose Function-word De-Attention (FDA) calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more aligned and robust VLMs. Specifically, we tested the FDA on 2 downstream tasks, 3 datasets, and 3 models, and evaluated all methods under 6 attacks. By comparing with existing SOTA defenses, our FDA shows superiority of FDA in boosting robustness.
Appendix
Appendix A Details for attacks and evaluation metrics
@@
Appendix B Full results for targeted attacks
@@
Flickr
$\nicefrac{{2}}{{255}}$
TeCoA 91.20 0.20 (+0.20) 5.30 (+4.90) 18.70 (+18.70) 20.40 (+20.00) 59.90 (+59.90) 86.40 (+86.00)
@@
Flickr
$\nicefrac{{2}}{{255}}$
TeCoA 81.44 0.10 (+0.08) 1.00 (+1.00) 16.42 (+16.40) 17.40 (+17.34) 56.90 (+56.90) 65.70 (+65.70)
@@
Appendix C Full results for untargeted attacks

### v3

@@ 4.1 Main Results
T2IR/I2TR. Results of T2IR/I2TR on ALBEF, TCL, and BLIP for all methods are given in Table.1. Overall on Flickr, our FDA consistently exhibits the best robustness with the best clean performance on all models over all other baselines for $\epsilon=\frac{2}{255}/\frac{4}{255}$, yielding a 22.26/14.69%, 14.29/13.55%, and 51.60/56.36% average ASR drop over all 3 attacks on ALBEF/TCL/BLIP, with a negligible 0.30/0.10%, 0.50/0.22%, and 0.70/0.46% performance drops in R@1 for T2IR/I2TR on each model, respectively. On MSCOCO, similar patterns exist as our FDA boosts the ASR drop by 9/14% for $\epsilon=\frac{2}{255}/\frac{4}{255}$ with a 0.1% clean performance boost on T2IR and a 0.07% drop on I2TR.
i. Attack-wise, on Flickr, our FDA exhibits the best defense against PGD and APGD in 22 out of 24 results, leading TeCoA/FARE by 60/65% on the BLIP model, demonstrating the effectiveness of FDA in enhancing robustness against various attacks. For the strongest adaptive attack, MAPGD, our FDA maintains its lead over TeCoA and FARE on ABLEF and BLIP, with an average lead by over 10%. Although our FDA shows more vulnerability against APGD on the TCL model, it retains the best comprehensive robustness of the other two baselines, yielding a 10-20% lead for $\epsilon=2/4$. TeCoA/FARE become ineffective for attacks with $\epsilon=4$, while our FDA retains its effectiveness facing stronger attacks. Similar trends also exist on MSCOCO.
ii. Performance-wise, all baseline methods (i.e., TeCoA and FARE) suffered from a performance drop by an average 4/3/9% on ALBEF/TCL/BLIP. Nevertheless, our FDA only causes minor or little drops of less than 1% for all models, yielding a lead of TeCoA and FARE by approximately 4%, 3%, and 7% on average, demonstrating the feasibility of paying less attention to function words for free robustness.
@@
4.2 Untargeted Attacks
We further evaluate the robustness against untargeted attacks. Thus, we retrained all models using TeCoA/FARE and their combination with our FDA to validate the effectiveness of FDA in defending against untargeted attacks.
For T2IR/I2TR, as presented in Table.3. Overall, FDA consistently boosts the robustness of TeCoA and FARE for all untargeted attacks on all models. Specifically, the scalability of FDA also persists with TeCoA/FARE: both methods benefit more from FDA on the larger backbone of BLIP, yielding a 4/3% robustness boost. Furthermore, we notice that FDA also boosts the clean performance of both methods on ALBEF considerably. For VG, we observe identical patterns: implementing FDA yields a solid robustness gain. For example, FARE experiences a significant robustness enhancement regarding untargeted attacks by 5%. In sum, our FDA compatibility works with both TeCoA and FARE to further boost their robustness against untargeted attacks.
Table 3: Robustness evaluations on ALBEF using FDA as a plug-and-play tool with TeCoA and FARE against untargeted attacks for Text-to-Image/Image-to-Text Retrieval. Results are averaged over T2IR and I2TR. Full results are provided in Sec.D of the Appendix. FDA consistently boosts clean performance and/or robustness against all attacks.
@@ 4.2 Untargeted Attacks
T2IR I2TR PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
@@
Table 4: Robustness evaluations of FDA as a plug-and-play tool with TeCoA and FARE against untargeted attacks for Visual Grounding. Results are averaged over Test-A and Test-B. Full results are provided in Sec.D of the Appendix. FDA consistently boosts clean performance and robustness against all attacks.
@@ 4.3 Ablation Study
All results are conducted on targeted attacks, hyperparameters of FDA: encoder, dictionary, layer/head, variant of FDA, and zero-shot performance. (See full results in Sec.E of the Appendix.)
@@
Table 6: Ablation studies on the layer/head index $L/H$ of FDA on Text-to-Image/Image-to-Text Retrieval on ALBEF, TCL and BLIP. Results are averaged over T2IR/I2TR. Shallower layers/heads (smaller $L/H$) consistently outperform over others on retrieval tasks.
@@
Table 7: Ablation studies on the layer/head index $L/H$ of FDA on VG on ALBEF, TCL, and BLIP. Results are given in average.
@@
Hyperparameters. The implementation of FDA, especially the macro-hyperparameters influencing where to implement, would largely impact the subsequent performance of models. We first provide relative ablation studies to help understand the mechanics and design of our FDA.
(1) Encoders&Dictionary: We start by comparing three implementations: FDA on text encoders, fusion encoders, and both, denoted as $\mathcal{T}$, $\mathcal{H}$, and $\mathcal{T}\&\mathcal{H}$. As presented in the top rows of Table.5, we find that $\mathcal{T}$ performs the worst among all, indicating that an early subtraction is insufficient for removing such subtraction. Although $\mathcal{T}\&\mathcal{H}$ provides a significant robustness boost, it costs an evident 2% performance drop on performance, implying that subtraction on both encoders is too strong and potentially causes contextual distortion. $\mathcal{H}$ performs the best as it helps models concentrate while preserving the contextual meaning.
For the dictionary, we use the off-the-shelf stopwords dictionary in (Li et al., 2020), containing 208 words/symbols, denoted as Full Dict. Furthermore, we use a shortlisted dictionary, by only using the most commonly used function words, containing 91 crucial function words, denoted as Shortlisted Dict. Both dictionary settings are trained with FDA $L^{all}$ to maximize their impacts on training. As presented in the lower row of Table.5, there are no significant performance gaps between the two settings, with Full Dict performing slightly worse regarding both clean and adversarial examples. We attribute the minor degradation to the length of the stopwords dictionary, which could unnecessarily skim words and distort the context. We provide the shortlisted dictionary in Sec.F of the Appendix.
[…]
FDA v.s. Masking & Other DA. We start by providing comparisons of our FDA and fine-tuning models by directly masking function words, including content words and nouns for sanity checks.
Table 8: Comparison for directly removing content words (CONT), nouns (NOUN), function words (FUNC), Determiner DA (DDA), and Adjective DA (ADA). $\Delta_{ASR}$ is presented using the average results for PGD and APGD.FDA shows significant advantages over directly masking and other de-attention variants.
Maksed Clean (R@1) ($\uparrow$) Avg ASR Drop
Words T2IR I2TR $\Delta_{ASR}$ ($\uparrow$)
N/A 95.90 85.60 -
CONT 21.50 11.10 -
NOUN 68.60 44.62 -
FUNC 94.00 80.86 $\uparrow$  1.56
DDA 95.60 85.42 $\uparrow$  9.28
ADA 95.50 85.38 $\uparrow$ 15.10
FDA 95.60 85.50 $\uparrow$ 23.07
For variants of FDA, we further compare our FDA with Adjective DA (ADA) and Determiner DA (DDA). Specifically, we choose determiners (DET) and adjectives because DET indicates using a small subset (i.e., a/an/the) of function words, while ADJ adopts a completely different set of words. Results are presented in Table.8. Note: We only test on PGD and APGD since MAPGD is not applicable for nouns and content words.
Overall, masking content words and nouns yields the largest performance drop, making it inviable for robustness evaluation. This aligns with the intuition that these words carry extensive semantic information crucial for VLM tasks. Furthermore, naively masking function words leads to evident performance drop ($\sim$1%) and brings negligible robustness. As a subset of FDA, DDA has identical clean performance but much less robustness. ADA also shows subpar performance compared with FDA. In sum, FDA leads the clean and adversarial performance among other variants.
@@
Zero-shot Performance. We also adopt three different head/layer settings of FDA for zero-shot T2IR/I2TR/VG on ALBEF and BLIP with $H^{0-5}$. Results are presented in Table.9. We find that $L^{all}$ performs the best for all tasks and all models. This not only suggests that $L^{all}$ serves as the most generalizable setting for multiple VL tasks and models, but also implies the feasibility of FDA for performance boost on zero-shot tasks.
Takeaway: Implementing FDA on the fusion encoder would be an ideal practice for all models. While we do not rule out potentially better performance by using different LL/$H$, one can adopt the setting in the paper on similar downstream tasks for stability, or simply use $L^{all},H^{0-5}$ as a safe choice.
Figure 3: Left: T-SNE of the vision-language embedding of vanilla VLM, FDA, FARE, and TeCoA. Our FDA is the most aligned model. Right: Comparison of text-image similarity for vanilla VLM versus VLM + FDA. Our FDA yields tightened alignment with larger similarities and smaller variances.
@@ 4.4 Analysis and Visualization
We notice that the results of APGD and MAPGD somewhat worsen after adversarial fine-tuning, e.g., TeCoA and FARE on ALBEF, FARE on BLIP in Table.1, etc. As previously illustrated in Fig.2, defending against targeted attacks requires a more aligned vision-language embedding. Consequently, we hypothesize that such abnormality potentially originates from the disruption in vision-language alignment brought by adversarial noise for enhanced robustness.
@@
4.5 Limitation
Although direct subtraction introduces superior efficiency, FDA could be potentially improved through a modular or algorithmic approach for more refined removal. Furthermore, we did not implement FDA to fine-tune a larger VLM or verify the effectiveness of FDA using PEFT methods such as LoRa (Hu et al., 2022) due to the hardware limitation. Additionally, our FDA is designed and implemented on backbones featuring a fusion encoder, but not been applied to projector-based models such as LLaVA (Liu et al., 2023) series and Qwen(Bai et al., 2023; Team and others, 2024; Bai et al., 2025; Yang et al., 2025) series, etc. Nevertheless, FDA could be seamlessly adopted on these backbones, which is a valuable exploration for future work. Finally, although we have built a shortlisted dictionary based on existing ones, it is possible that the optimal dictionary might have less/more words, which requires enormous effort for validation.
@@ 5 Conclusion
In this paper, we propose Function-word De-Attention (FDA) calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more aligned and robust VLMs. Specifically, we tested the FDA on 2 downstream tasks, 3 datasets, and 3 models, and evaluated all methods under 6 attacks. By comparing with existing SOTA defenses, our FDA shows superiority of FDA in boosting robustness and clean performance. We also provide an in-depth analysis of FDA and validate its boost on zero-shot performance. Our work validate the redundancy of function words in vision-language models and the subsequent vulnerability brought by these words, calling for more research interest in further exploration.
Acknowledgement
This work was supported by National Key Research and Development Program of China (2023YFB3107401), the National Natural Science Foundation of China (T2341003, 62521002, U2441240, U24B20185, 62376210, 62132011, 62406240, U244120060).
Ethics Statement
We acknowledge that all authors of our papers are required to read the Code of Ethics, adhere to it, and explicitly acknowledge this during the submission process. Contribute to society and to human well-being. All authors: i) uphold high standards of scientific excellence; ii) avoid harm; iii) be honest, trustworthy, and transparent; iv) be fair and take action to avoid discrimination; v) respect the work required to produce new ideas and artefacts; vi) respect privacy; vii) honour confidentiality.
Reproducibility Statement
The detailed information about the implementation of FDA is provided in Section.4. The data used in this paper is open-source, and the details/full results are stated in the Appendix. The code and the checkpoint will be publicly available along with sufficient instructions to faithfully reproduce the main experimental results/visualization, and detailed instructions to transfer to other backbones.
Appendix for Pay Less Attention to Function Words for Free Robustness of Vision-Language Models
Appendix A FDA v.s. Adaptive Words selection
[…]
Appendix B Details for attacks and evaluation metrics
@@
in Dictionary T2IR I2TR $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ Avg
SIM-N 25.95 95.90 85.50 2.44 12.81 $\uparrow$  7.62
SIM-$2\delta$ 74.53 95.60 85.32 9.38 13.86 $\uparrow$  8.39
SIM-$\delta$ 79.49 95.30 85.38 12.85 11.54 $\uparrow$ 12.41
FDA 100.00 95.90 85.50 27.61 18.53 $\uparrow$ 23.07
Appendix C Full results for targeted attacks
@@
Flickr $\nicefrac{{2}}{{255}}$ TeCoA 91.20 0.20 (+0.20) 5.30 (+4.90) 18.70 (+18.70) 20.40 (+20.00) 59.90 (+59.90) 86.40 (+86.00)
@@
Flickr $\nicefrac{{2}}{{255}}$ TeCoA 81.44 0.10 (+0.08) 1.00 (+1.00) 16.42 (+16.40) 17.40 (+17.34) 56.90 (+56.90) 65.70 (+65.70)
@@
Appendix D Full results for untargeted attacks